In a lottery game, a player picks 8 numbers from 1 to 52. If 7 of the 8 numbers match those drawn, the player wins second prize. What is the probability of winning this prize? Be sure to leave your answer as a fraction in order to earn credit.
Answer #1

Total numbers are 1 to 52.

Players picks up 8 numbers which can be done in 52C8 ways

For winning 7 numbers need to be from the winning group of 8 numbers.

No of ways in which these 7 numbers can be picked = 8C7 ways

P(players getting a winning set) = total number of favourable outcomes / no of outcomes

P(players getting a winning set) = 8C7 / 52C8

P(players getting a winning set) = 8 / 752538150 = 4/ 376269075